OpenStudy (anonymous):

Please help. A cube has sides of length √5. I need to find the area of one side of the cube. Find the volume of the cube. find the length of the diagonal of the sides.

OpenStudy (anonymous):

area is radical 5 times radical 5

OpenStudy (anonymous):

the volume is \[\left( \sqrt{5} \right)^{3}\]

OpenStudy (anonymous):

and to find the diagonal line you have to use the pythagorean theorem

OpenStudy (anonymous):

that diagonal line is the hypoteneus and the radical 5's are the two sides thus that...

OpenStudy (anonymous):

How would I find the distance from A to B in a figure they show in the textbook?

OpenStudy (anonymous):

the diagonal line?

OpenStudy (anonymous):

yes, it is a digagonal line from one corner to the other corner

OpenStudy (anonymous):

since its a cube...we can assume that one side of the cube is a perfect 2 dimentional square...so each side is \[\sqrt{5}\]

OpenStudy (anonymous):

you notice that each side of the square of one side of the cube is measured the same thus all the sides are radical 5

OpenStudy (anonymous):

use the pythagorum theorem to find what the diagonal line's length is

OpenStudy (anonymous):

you are trying to find \[c ^{2}; \in the formula a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

you have a^2 and b^2 already

OpenStudy (anonymous):

How do I do that when there are no numbers provided. It is just A to B.

OpenStudy (anonymous):

well it tells you in the problem that the side of the square is \[\sqrt{5}\]

OpenStudy (anonymous):

this applies to all the side so the cube

OpenStudy (anonymous):

So to figure it out I would use √5²+√5²=c²

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you very much. Hopefully I can figure this out.

OpenStudy (anonymous):

i wish you the best of luck :)